The floor of a shed has an area of 88 square feet. The floor is in the shape of a rectangle whose length is 5 feet less than twice the width. Find the length and the width of the floor of the shed. Use the formula, area=length*width.
Answer by amarjeeth123(574)

Let the width be x.
Then the length is (2x-5).
area=length*width
x(2x-5)=88
2x^2-5x-88=0
2x^2-16x+11x-88=0
2x(x-8)+11(x-8)=0
(2x+11)(x-8)=0
x=8 is the only positive solution.
The width of the shed is 8 feet.
The length of the shed is 11 feet.